# Algorithmic Misjudgement in Google Search Results: Evidence from Auditing the US Online Electoral Information Environment

### v1

@@ 5.2.1. Domains
Table 4 shows the 10 non-federal government domains which make up the highest proportion of organic results, and together compose 35.96% of appearances for non-federal domains, or 11.8% of organic results overall. Because the number of locations for which data was collected varied per state, it is not appropriate to directly compare the frequency with which these domains occur to each other. However, it is clear that these domains suffer from high rates of mistargeting, with many appearing more often in other states than the states for which they serve. Notably, wakegov.com, a county government site for Wake County, North Carolina, and votebrevard.gov, the Supervisor of Elections site for Brevard County, Florida, are mistargeted over 94% of the time. Overall, 84 domains were mistargeted more often than they were correctly targeted, and they compose 44% of non-federal organic results (14.45% of all organic results). In contrast, 245 domains are correctly targeted more often than they are mistargeted, and they compose 41% of non-federal organic results (13.5% of all organic results).
Thus, while the majority of non-federal government sources are always correctly targeted at the state-level, these sources make up only a minority of all organic results; a relatively small subset of domains are mistargeted more often than they are correctly targeted, and these mistargeted results show frequently among organic results.
5.2.2. Misjudgement Levels by Query Category and State
Table 5 shows the percent of all organic results that are correctly targeted to the appropriate locality, and Table 6 shows the mistargeting rates for each state, county, and local government sources for different query categories. 71% of all appearances of locality-specific government websites in organic results were mistargeted, with the proportion of mistargeting highest for Candidates and Issues queries. Voting Logistics queries, which see the highest proportion of government results by far, have the lowest mistargeting rate of the 3 query categories, albeit still 69.4%.
@@ 5.2.2. Misjudgement Levels by Query Category and State
Table 5. Percent of all organic results that are correctly targeted to the appropriate locality for state, county, and local sources. While 32.89% of all organic results can be attributed to locality-specific government sources, only 9.47% of organic results are composed of correctly targeted government sources.
@@
Table 6. Proportion of government results that were mistargeted; the denominator here is the total appearances of a source type, rather than all organic results. Overall, 71% of locality-specific government organic results were mistargeted, with the mistargeting rate being highest for county government sources at nearly 92%.
Across all queries, county-level government sources suffer from the highest levels of mistargeting, with 91.9% of these results mistargted. State-level government domains, which appear the most among organic results, have the lowest mistargeting rate of the locality-specific sources at about 58.3%. Mistargeting of state-level government sources is lowest for Voting Logistics queries at 51.75%, notable since state-level sources are the most common government source in this query category. Lastly, local government sources suffer from a mistargeting rate of 76%, unfortunate considering the minimal appearances of local government sources in organic results.
There is some variation in the extent to which government domains appear in different states, as well as mistargeting rates by state, as visualized in Figure 2. Studying this variation is left to future work.
Figure 2. Variation in the extent to which government sources appear in different states. much of the mistargeting for local government sources (and county government sources, to a lesser extent) can be attributed to other government sources within the respective state, indicating a lack of precision with targeting locality-specific government domains to the appropriate audience when relying only on geocoordinates to represent user location.
@@ 5.2.3. Mistargeting Severity
With respect to rank, while only 28.8% of non-federal government sources are correctly targeted, correctly targeted results generally appear in higher ranks among organic results than mistargeted results (Figure 3). The median rank of a correctly-targeted government result was 3, while the median rank of a mistargeted government result was 7. Moreover, for the 43.7% of SERPs that contained both correctly-targeted and mistargeted government results, the first correctly-targeted result appeared above the first mistargeted result 83.9% of the time. A mistargeted government result appeared in the first organic position for only 3.3% of SERPs analyzed, while a correctly-targeted government result appeared in the first organic position for 19.3% of SERPs. Given users tend to focus on search results in the top positions (Pan et al., 2007), these findings suggest that while mistargeting rates are high, their severity, within the context of impact on a user, may be low. Additionally, these findings may indicate that the search algorithm prioritizes non-federal government sources in ranked results only when there is a high likelihood that the source geoprovenance matches the user’s locations; otherwise, non-federal government sources appear in lower ranked positions, and other content (such as a news or non-profit organization) is given the top ranks.
@@
Figure 3. While only 28.8% of non-federal government domains are correctly targeted, these results usually appear in higher ranks among organic results. On the 43.7% of SERPs that have both correctly-targeted and mistargeted results, the correctly targeted result appeared above the mistargeted result 83.9% of the time.
@@
With respect to the nature of mistargeted results, a high percentage of mistargeting for local government sources can be attributed to local government sources for places in the same state as the search location. While the overall mistargeting rate for local government sources is 76%, 64% of local government results are for other localities in the correct state. Figure 2 visualizes this phenomenon; in most states, nearly all of the mistargeted local government results are from other local government sources in the respective state. A similar phenomenon with targeting county-level government sources is observed, though to a lesser degree and with more variation between states. These findings imply a lack of precision with targeting locality-specific government sources to the appropriate audience when using general queries that do not specify location information. While a similar lack of precision has been observed with geotargeting of Ads on Google (Bandy and Hecht, 2021), the consequences of this lack of precision is greater in the context of “low information voters” [^20] https://en.wikipedia.org/wiki/Low_information_voter searching for election information as it relates to their locale, given that other county or locality information is likely not be applicable to them.
Lastly, while we find no correlation between the number of local governments and the percent that local government sources compose in organic results, we find a moderate, positive correlation between the number of local governments in a state and the mistargeting rate for local government sources (r=0.36), and for the correlation between number of local governments and the rate of in-state mistargeting (r=0.37), suggesting to some degree that a greater number of local governments is associated with less precise targeting of local government sources in organic search.
5.3. Presence of Election Information Boxes
Between Oct 18-24, only 39 queries showed an electoral box (across most locations), though this number jumped to 73 on Oct 25 (two weeks before election) and went up to 83 on election day. Notably, on election day, all locations had at least one election box. Once the elections were over, the number of queries with an election box went back in the 30s, but remained consistent across most locations.
These results indicate that the presence of election information boxes is not location-specific, but depends on the date and query phrase. Only 86 out of 161 query phrases (53%) had at least one box on at least one date. Furthermore, 66% of these queries (57 out of 86) belonged to the list of queries we collected from Google Trends. These queries made up 77% of all queries from that source (57 out of 74). This suggests that Google might be showing election information boxes for popular query phrases of a certain type. However, a more detailed analysis (with a more expansive set of queries) might be needed to understand what triggers these boxes.
6. Discussion and Future Work
6.1. How Algorithmic Misjudgement Shapes the Electoral Information Environment
The results of our first research question reveal that about 40.6% of all websites are government sources and that combined they contribute 39.97% of organic results. This might look like an excellent outcome, given the almost perfect matching between the proportion of domains and the proportion of their occurrences in the search results. However, our deeper analysis of the results reveals several issues that are hidden, when one only considers proportions.
@@ 6.1. How Algorithmic Misjudgement Shapes the Electoral Information Environment
For one, the top ten non-federal government sources show up disproportionally often, contributing to 11.8% of all organic results (see Table 4). As a reference point, all 1,848 non-federal sources combined make up only 32.89% of all results (i.e., about 1/3 of their combined results are accounted for by these ten domains). This phenomenon is reminiscent of source concentration and mainstreaming effects found in previous work (Trielli and Diakopoulos, 2022, 2019) for news sources. While it might not be harmful to an individual to read national news from New York Times more than from Baltimore Sun (although, collectively, source concentration has led to the so-called ’news desert’ phenomenon (Abernathy, 2018)), for elections, all out-of-state web domains might be irrelevant to a voter, when it comes to their voting information. Our results make it clear that Google Search is treating electoral information needs as any other type of information and is populating the organic search results with popular websites, instead of locally-relevant websites.
Further, while correctly-targeted government sources are usually ranked higher than those that are mistargeted (Figure 3), indicating that Google is capable of getting things right (19.3% of SERPs have a correctly targeted government website at the first position, mostly a state government), 71.19% of government website occurrences were mistargeted. Overall, our analysis indicates a lack of consistency in the search algorithm’s treatment of locality-specific government sources, leaving most government domains, especially those of local governments, appearing minimally, inconsistently, or missing entirely from election-related search results. We consider the case of mistargeted websites and that of the absence of locality-specific websites as instances of algorithmic misjudgement that need to be corrected.
Google Search is aware of the importance of authoritative information about elections, as indicated by the consistent inclusion of “election information panels” in prominent SERP positions across locations. Google should consider that showing locality-specific web domains in the organic results might contribute to strengthening the civic infrastructure of the election by increasing the visibility of information from elected officials.
@@ 6.2. Limitations
There are limitations to our study. First, we focus exclusively on government web domains in search results, but in reality there are other types of information that users encounter in search results, including results from other types of sources and curated information panels. Additionally, our audit was conducted when there were around 10 organic results per SERP, while since December 2022 (after our data collection), Google has moved to “continous scrolling”, which loads up to 60 results at once. [^21] https://www.theverge.com/2022/12/6/23495973/google-search-desktop-continuous-scrolling Second, we rely on geocoordinates to serve as a proxy for locations of real users and do not study how including location information in the search query (eg. “voting near me Boston, MA”) impacts the composition and mistargeting of government search results. While the queries come from Google Trends and real users, we do not study how users actually interact with government sources in search results. Moreover, differences in popularity of search terms between those that come from Google Trends and those from real users may impact the composition of search results (i.e. more popular queries may have better results).Third, we analyze data from 786 locations across 25 states, a size which is sufficient for statistical analysis but does not capture all of the United States. Finally, this study only focuses on the United States, given that electoral laws and election authorities are different state by state, making this kind of analysis meaningful. In other countries, where there is a single election authority, the needs and assumptions might be different.
@@ Acknowledgements.
We are grateful to members of our lab and funding from XXX.

### v2

@@ 5.2.1. Domains
Table 4 shows the 10 non-federal government domains which make up the highest proportion of organic results, and together compose 35.96% of appearances for non-federal domains, or 11.8% of organic results overall. Because the number of locations for which data was collected varied per state, it is not appropriate to directly compare the frequency with which these domains occur to each other. However, it is clear that these domains suffer from high rates of mistargeting, with many appearing more often in other states than the states for which they serve. Notably, wakegov.com, a county government site for Wake County, North Carolina, and votebrevard.gov, the Supervisor of Elections site for Brevard County, Florida, are mistargeted over 94% of the time. Overall, 84 domains were mistargeted more often than they were correctly targeted, and they compose 44% of non-federal organic results (14.45% of all organic results). In contrast, 245 domains are correctly targeted more often than they are mistargeted, and they compose 41% of non-federal organic results (13.5% of all organic results). Thus, while the majority of non-federal government sources are always correctly targeted at the state-level, these sources make up only a minority of all organic results; a relatively small subset of domains are mistargeted more often than they are correctly targeted, and these mistargeted results show frequently among organic results.
@@
Table 5. Percent of all organic results that are correctly targeted to the appropriate locality for state, county, and local sources; the denominator here is the number of all organic results (about 31 million). While 32.89% of all organic results can be attributed to locality-specific government sources, only 9.47% of organic results are composed of correctly targeted government sources.
5.2.2. Mistargeting by Query Category and State
Table 5 shows the percent of all organic results that are correctly targeted to the appropriate locality, and Table 6 shows the mistargeting rates for each state, county, and local government sources for different query categories. Across all SERPs, 71% of all appearances of locality-specific government websites in organic results were mistargeted. Mistargeting is highest among Candidates and Issues queries at 78.61%. Voting Logistics queries, which see the highest proportion of government results by far, have the lowest mistargeting rate of the 3 query categories, albeit still 69.4%.
@@ 5.2.2. Mistargeting by Query Category and State
Table 6. Percent of government results that were mistargeted. Mistargeting rate is calculated by subtracting the proportion of correctly targeted (Table 5) from the total proportion (Table 3), and dividing by the total proportion. . Overall, 71% of locality-specific government organic results were mistargeted, with the mistargeting rate being highest for county government sources at nearly 92%.
Across all queries, county-level government sources suffer from the highest levels of mistargeting, at 91.9%. State-level government domains have the lowest mistargeting rate of the locality-specific sources at about 58.3%. Mistargeting of state-level government sources is lowest for Voting Logistics queries at 51.75%, notable since state-level sources are the most commonly appearing government source in this query category. Lastly, local government sources suffer from a mistargeting rate of 76%, unfortunate considering the minimal appearances of local government sources in organic results.
Figure 2. While only 28.8% of non-federal government domains are correctly targeted, these results usually appear in higher ranks among organic results. On the 43.7% of SERPs that have both correctly-targeted and mistargeted results, the correctly targeted result appeared above the mistargeted result 83.9% of the time. Thus, while mistargeting rates are high, their severity, within the context of impact on a user, may be low, given that users often rely on top ranking results.
Figure 3. Much of the mistargeting of local government sources (and county government sources, to a lesser extent) can be explained by the appearance of sources that serve other localities in the same state as the search location. This indicates a lack of precision with targeting locality-specific government domains to their specific audience when relying only on geocoordinates to represent user location.
@@ 5.2.3. Mistargeting Severity
With respect to rank, while only 28.8% of non-federal government sources are correctly targeted, these correctly targeted results generally appear in higher ranks among organic results than mistargeted results (Figure 2). The median rank of a correctly-targeted government result was 3, while the median rank of a mistargeted government result was 7. At the SERP level, 5.6% of SERPs analyzed had only correctly targeted results, 19.9% of SERPs had only mistargeted results, 43.7% of SERPs had both correctly-targeted and mistargeted government results, and the remaining 30.7% of SERPs had no subnational government source. Thus, just under half of SERPs (49.3%) had at least one correctly targeted result. Moreover, for the 43.7% of SERPs that contained both correctly-targeted and mistargeted government results, the first correctly-targeted result appeared above the first mistargeted result 83.9% of the time. Lastly, a mistargeted government result appeared in the first organic position for only 3.3% of SERPs analyzed, while a correctly-targeted government result appeared in the first organic position for 19.3% of SERPs.
Given that users tend to focus on search results in the top positions (Pan et al., 2007), these findings suggest that while mistargeting rates are high, their severity, within the context of impact on a user, may be low. Additionally, these findings may indicate that the search algorithm prioritizes non-federal government sources in ranked results only when there is a high likelihood that the source geoprovenance matches the user’s locations; otherwise, non-federal government sources appear in lower ranked positions, and other content (such as a news or non-profit organizations) is given the top ranks.
With respect to the nature of mistargeted results, a high percentage of mistargeting for local government sources can be explained by the appearance of sources that serve other localities in the same state as the search location. While the overall mistargeting rate for local government sources is 76%, 64% of local government results are for other localities in the correct state. Figure 3 visualizes this phenomenon; in most states, nearly all of the mistargeted local government results are from other local government sources in the respective state.
@@
A similar phenomenon with targeting county-level government sources is observed, though to a lesser degree and with more variation between states. These findings imply a lack of precision with targeting locality-specific government sources to the appropriate audience when using general queries that do not specify location information. While a similar lack of precision has been observed with geotargeting of Ads on Google (Bandy and Hecht, 2021), the consequences of this lack of precision is greater in the context of “low information voters” [^19] https://en.wikipedia.org/wiki/Low_information_voter searching for election information as it relates to their locale, given that other county or locality information is likely not be applicable to them.
6. Discussion, Limitations, and Future Work
6.1. Discussion
This paper studies the extent to which local, county, and state government sources appear in election-related search results, as well as the extent to which these sources are correctly targeted to the locations which they serve. We label misalignment between the geographic area that non-federal domains serve and the locations for which they appear in search results as mistargeting, a subtype of algorithmic misjudgement in which the search algorithm targets locality-specific information to users in different, incorrect locations.
6.1.1. Overrepresentation of Some Locality-Specific Government Sources
@@
With respect to our first research question, we find that 40.6% (1,848/4,556) of websites that appeared in organic results are government sources, and combined, these sources compose 39.97% of all organic results. This might look like an excellent outcome, given the almost perfect matching between the proportion of domains and the proportion of their occurrences in the search results. However, we find that that some locality-specific government domains are over represented and frequently mistargeted in organic results. About 1/3 of all occurrences of non-federal government sources can be attributed to ten domains, with these domains frequently appearing in locations for which they do not serve (see Table 4). This phenomenon is reminiscent of source concentration and mainstreaming effects found in previous work for news sources (Trielli and Diakopoulos, 2022, 2019). While it might not be harmful to an individual to read national news from New York Times more than from Baltimore Sun (although, collectively, source concentration has led to the so-called ’news desert’ phenomenon (Abernathy, 2018)), for elections, all out-of-state web domains might be irrelevant to a voter, when it comes to their voting information.
6.1.2. Extensive Mistargeting in Low Ranks
Furthermore, we find extensive mistargeting of locality-specific government sources: 71.18% of all occurrences of non-federal government sources were mistargeted, with county-level government sources having the highest mistargeting rate at 91.94% (Table 6). However, we also find that correctly targeted results generally appear in higher ranks among organic results than mistargeted results (Figure 2), and for the 43.7% of SERPs that contained both correctly targeted and mistargeted government results, the first correctly targeted result appeared above the first mistargeted result 83.9% of the time. In the absence of user behavior data, these findings are encouraging that the severity of harm to a user due to mistargeted results may be low, since low-ranking mistargeted results are usually accompanied by high-ranking correctly targeted results.
6.1.3. Government Sources, SERP Composition, and Consequences of Mistargeting
The study of mistargeted government results in election SERPs is motivated by how voters and election officials position election officials and government sources as important, trusted conveyors of election information. [^20] See Introduction; https://bipartisanpolicy.org/blog/new-survey-data-election-information/ and https://www.nass.org/initiatives/trustedinfo It is not our aim to claim that government sources are the most (or only) relevant type of source for electoral information. Indeed, it is conceivable that higher quality, non-government sources containing relevant information may be more beneficial to a user than lower quality, correctly-targeted government sources. Exploring the extent to which government sources are quality sources for electoral information is beyond the scope of this paper. However, regardless of whether government sources should be considered as the best and highest quality, our audit shows not only that locality-specific government sources are showing in SERPs for different types of queries (even if minimally), but also that these results are often not showing up in the correct places. Questioning whether a locality-specific government source is correctly targeted to its constituents perhaps precedes questions about whether this source type is most favorable for a user. At best, mistargeting of government sources among organic search results constitutes wasted space on a SERP that could otherwise be used to provide voters with relevant electoral information. At worst, mistargeting constitutes a failure in Google’s ability to facilitate (easy) access to accurate, authoritative, and locality-specific electoral information to voters who turn to the search engine with their election-related information needs.
Google states that it is committed to providing voters with authoritative information about US elections. Our focus on mistargeting of government sources among organic results reveals room for improvement in this regard. Showing locality-specific government sources in organic results and ensuring that these sources are correctly targeted to their respective constituents are necessary criteria for ensuring that Google Search can appropriately provide users across different locations access to authoritative and locally-relevant election information.
@@ 6.2. Limitations
There are limitations to our study. First, we focus exclusively on government web domains in search results, but in reality there are other types of information that users encounter in search results, including results from other types of sources and curated information panels on the SERP. During the 2022 US midterm elections in particular, Google populated election-related SERPs with election information panels, many of which contained curated, state-specific information; an analysis of these panels is necessary to get a complete understanding of how locality-specific electoral information is surfaced among search results, but this is left to future work. Additionally, our audit was conducted when there were around 10 organic results per SERP, though since December 2022 (after our data collection), Google has moved to “continuous scrolling”, which loads up to 60 results at once. [^21] https://www.theverge.com/2022/12/6/23495973/google-search-desktop-continuous-scrolling
Second, we rely on geocoordinates to serve as a proxy for locations of real users and do not study how including location information in the search query (e.g. “voting near me Boston, MA”) impacts the composition and mistargeting of government search results. We also do not study whether targeting based on geocoordinates differs from IP-based targeting, nor do we consider how user profiles impact targeting. It is conceivable that IP-based targeting and/or user profiles lead to more location precision and thus less mistargeting among organic results. Furthermore, while the queries come from Google Trends and real users, we do not study how users actually interact with government sources in search results.
We analyze data from 786 locations across 25 states, a size which is sufficient for statistical analysis but does not capture all of the United States. Additionally, we do not analyze how mistargeting varies by search location features (including the extent to which the congressional election for a location is competitive) nor by date. Finally, this study only focuses on the United States, given that electoral laws and election authorities are different state by state, making this kind of analysis meaningful. In other countries, where there is a single election authority, the needs and assumptions might be different. Indeed, it is unclear if these results can be generalized to other countries with similar political systems to the United States. In addition to the similarity of political systems, one would have to consider the strength of a country’s online government and electoral information infrastructure when generalizing results.
@@ Acknowledgements.
We are grateful to members of Wellesley Cred Lab and acknowledge funding from NSF grant ISS 1751087.